A department store sells a pair of shoes for $73 and an assortment
of bags for $44 each. In one day, the store sells n pairs of shoes
and (n - 7) bags. Which expressions represent the day’s total sales
for shoes and bags at the store?
A) 73n + 44n + 308
B) 44(n - 7) + 73n
C) 44n + 73(n - 7)
D) 73n + 44(n - 7)
E) n(73 + 44) - 7
F) 117n - 308
G) 117n - 7

Found 3 solutions by Theo, ikleyn, greenestamps:
Answer by Theo(13342)

i get 73n + 44(n-7).

the store gets 73 dollars a pair of shoes and 44 dollars a bag.

in one day, the store sells n pairs of shoes and (n-7) bags.

n pairs of shows * 73 dollars a pair = 73 * n. = 73n.
(n-7) bags * 44 dollars a bag = 44 * (n - 7) = 44(n-7).
the total revenue is 73 * n + 44 * (n - 7) = 73n + 44(n-7).
that's selection D.


Answer by ikleyn(52779)

.

        The complete answer is options  (D)  and  (F).

Indeed, the correct formula for the total sales is 

    73n + 44(n-7)


and all its equivalent forms.


(D) is literally 73n + 44(n-7).


But if you open parentheses, you will find other equivalent form

    73n + 44(n-7) = 73n + 44n - 308 = 117n - 308,

which is option (F).

Solved (correctly).



Answer by greenestamps(13200)



It seems it takes three responses to get the right answer....

Answers B and D are the same, and answer F is the equivalent of either B or D after parentheses are removed and like terms are combined.

ANSWERS: B, D, and F